Seven curliy fries and four plain fries cost $16.83. Five curlies and 12 plains cost $23.45. How much does one order of curlies cost? One order of plains?
Answer by checkley77(12844)

7c+4p=16.83 multiply this equation by -3 & add
5c+12p=23.45
-21c-12p=-50.49
-------------------------
-16c=27.04
c=-27.04/-16
c=1.69 for one curlie fries.
7*1.69+4p=16.83
11.83+4p=16.83
4p=16.83-11.83
4p=5.00
p=5.00/4
p=1.25 for one plain fries.
Proof:
5*1.69+12*1.25=23.45
8.45+15=23.45